OpenStudy (anonymous):

Q. Show that the function \(f(x)\) defined below attains a unique minimum for \(x > 0\). What is the minimum value of the function? What is the value of \(x\) at which the minimum is attained? $$\hspace{3cm} f(x) = x^2+x+\frac1x+\frac1{x^2}\quad \text{for}\;\;x\neq0$$ Is there an elegant way of proving this? Instead of differentiating twice and solving cubic equation.

OpenStudy (anonymous):

this looks simmilar to Laurent series

OpenStudy (blockcolder):

Rewrite the function like this: \[f(x)=x^2-2+\frac{1}{x^2}+x-2+\frac{1}{x}+4\\ =\left( x-\frac{1}{x}\right)^2+\left( \sqrt{x}-\frac{1}{\sqrt{x}}\right)^2+4.\]

OpenStudy (anonymous):

Is proving that some function \(g(x)\) defined as \(g(x) = x +\frac{1}{x}\quad\)for\(\;x>0\), is minimized when \(x=1\), sufficient to prove \(h(x) = u(x)+\frac{1}{u(x)}\quad\)for \(x>0\) is minimized when \(u(x)=1\)?

OpenStudy (blockcolder):

You don't need to see it that way. Look at it this way: The expression is composed of squares, which attain a minimum value of 0 at x=1.

OpenStudy (anonymous):

Ohh hmm sorry :/ Yeah, that's much better.

OpenStudy (anonymous):

Nice blockcoder. Thanks. I wonder why it didn't cross my mind :/